# The Importance of Being Discrete: Measuring the Impact of Discretization in End-to-End Differentially Private Synthetic Data

### v2

@@ 7 Conclusion
Determining the optimal number of bins is not a straightforward task. For instance, utility initially improves with more bins but then degrades with an excessive number of bins. Moreover, training multiple models to identify the best configuration can negatively impact privacy by depleting the privacy budget, besides requiring significant computational resources. This motivated us to derive an optimized formula to select the ideal number of bins, helping bypass the need for extensive model training.
@@ Appendix A Scikit-Learn Examples
In Figure 16 and 17, we present two examples, visually highlighting the importance of discretization in clustering and classification. (Inspired by Scikit-Learn, see https://scikit-learn.org/stable/auto_examples/preprocessing/plot_discretization_strategies.html as well as https://scikit-learn.org/stable/auto_examples/preprocessing/plot_discretization_classification.html).
In the first example, shown in Figure 16, we apply DP discretization to a clustering task on a toy dataset. The figure uses different colors to denote regions where the discretized encoding remains constant. It illustrates how different discretizers encode the data differently and, except for uniform, how the privacy budget influences the shape, size, and number of resulting bins. For instance, quantile does not perform particularly well, as it merges records from different clusters across all budgets. Unsurprisingly, k-means performs consistently well, while PrivTree reduces the number of bins as the privacy budget decreases.
The second example, visualized in Figure 17, compares the performance of logistic regression with and without DP discretization on three toy datasets. The shade of the colors reflects the confidence of the fitted classifiers, highlighting how discretization affects decision boundaries. As expected, discretization does not help in the nearly linearly separable case (third row). However, in the other two cases, all discretizers except one, at $\epsilon=\infty$, produce data representations that improve classifier accuracy. Notably, classifiers trained on discretized data with $\epsilon=1$ outperform the baseline for the circles data (second row). This improvement is likely due to the limitations of logistic regression; more advanced classifiers, like random forests, would not face these challenges initially. Nevertheless, this example shows that discretization can enhance modeling in certain scenarios. Overall, both quantile and k-means discretizers demonstrate strong performance, while PrivTree underperforms.

### v4

@@ 7 Conclusion
Determining the optimal number of bins is not a straightforward task. For instance, utility initially improves with more bins but then degrades with an excessive number of bins. Moreover, training multiple models to identify the best configuration can negatively impact privacy by depleting the privacy budget, besides requiring significant computational resources. This motivated us to optimize an existing method for selecting the ideal number of bins, helping bypass the need for extensive model training.
@@
Acknowledgments
We are grateful to the ACM CCS Program Committee for their valuable feedback, which helped us significantly improve our paper.
@@ Appendix A Scikit-Learn Examples
In Figure˜16 and 17, we present two examples, visually highlighting the importance of discretization in clustering and classification. (Inspired by Scikit-Learn, see https://scikit-learn.org/stable/auto_examples/preprocessing/plot_discretization_strategies.html as well as https://scikit-learn.org/stable/auto_examples/preprocessing/plot_discretization_classification.html).
In the first example, shown in Figure˜16, we apply DP discretization to a clustering task on a toy dataset. The figure uses different colors to denote regions where the discretized encoding remains constant. It illustrates how different discretizers encode the data differently and, except for uniform, how the privacy budget influences the shape, size, and number of resulting bins. For instance, quantile does not perform particularly well, as it merges records from different clusters across all budgets. Unsurprisingly, k-means performs consistently well, while PrivTree reduces the number of bins as the privacy budget decreases.
The second example, visualized in Figure˜17, compares the performance of logistic regression with and without DP discretization on three toy datasets. The shade of the colors reflects the confidence of the fitted classifiers, highlighting how discretization affects decision boundaries. As expected, discretization does not help in the nearly linearly separable case (third row). However, in the other two cases, all discretizers except one, at $\epsilon=\infty$, produce data representations that improve classifier accuracy. Notably, classifiers trained on discretized data with $\epsilon=1$ outperform the baseline for the circles data (second row). This improvement is likely due to the limitations of logistic regression; more advanced classifiers, like random forests, would not face these challenges initially. Nevertheless, this example shows that discretization can enhance modeling in certain scenarios. Overall, both quantile and k-means discretizers demonstrate strong performance, while PrivTree underperforms.
Appendix B RQ2: Discretization and DP Generative Models (Additional Experiment)
In Figure˜18, we present disaggregated metrics (namely, maximum percentile distance, discriminator similarity, and query similarity) of the four discretizers across a range of bins and two privacy budgets in US2. The overall behavior of the discretizers is consistent across the metrics, as discussed in Section˜5.2.
Appendix C RQ4: Discretization and DP Domain Extraction (Additional Experiment)
In Figure˜19, we present results from running Querybased [33] in PS1 on the Wine dataset with three data domain extraction strategies and two target records. The findings closely mirror those obtained with GroundHog [77], as discussed in Section˜5.4.